OpenStudy (anonymous):

Find all the real roots of the polynomial. x^3+x^2-27x-27 Why isn't the answer -1?

OpenStudy (anonymous):

it equals to zero so why is it wrong?

OpenStudy (slaaibak):

One of the roots is definitely -1. x^3 + x^2 - 27x - 27 -1 is a root, so it's divisible by (x+1) factoring: (x+1)(x^2 - 27) x^2 - 27 = 0 x= +-3sqrt3

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

use the rational roots test... r/s

OpenStudy (anonymous):

when u do, u get +-1, +-3,+-9,+-27

OpenStudy (anonymous):

and when u plug those into the function, the only one that equals to zero is -1

OpenStudy (slaaibak):

That will not give you all the roots. http://www.wolframalpha.com/input/?i=x^3%2Bx^2-27x-27 wolframalpha seems to agree with me. So I don't think I'm wrong.

OpenStudy (anonymous):

i put -+ \[\sqrt{3}\]

OpenStudy (anonymous):

i mean -+ 3 sqrt{3}

OpenStudy (anonymous):

and it said im wrong on my hw site

OpenStudy (slaaibak):

Then your hw site is wrong.

OpenStudy (anonymous):

doubt it

OpenStudy (slaaibak):

I don't. Or it's user error. wolframalpha > your hw site

OpenStudy (anonymous):

ok i just did a very similar problem the exact same way i did this one and got the right answer -_-

OpenStudy (anonymous):

i have no idea what im doing wrong for this one

OpenStudy (slaaibak):

maybe you have the wrong input to the site.

OpenStudy (shadowfiend):

The problem with using the rational roots test is that \(\sqrt{3}\) is not a rational root, it's an irrational one.

OpenStudy (shadowfiend):

Try putting in sqrt{3} and -sqrt{3} separately on the site, perhaps?

OpenStudy (anonymous):

thats not going to do it

OpenStudy (anonymous):

my hw website gives u the option to use the square root sign already... i use webassign btw

OpenStudy (anonymous):

by the way the answer was -1, sqrt{27}, and - sqrt{27}

OpenStudy (slaaibak):

sqrt27 = 3sqrt3

OpenStudy (anonymous):

i know but it didnt accept it the way you had it so i tried every other way until it accepted it

OpenStudy (shadowfiend):

Ouch. That's pretty retarded. I'm sorry to hear the system isn't smart enough to understand that :/